OpenStudy (anonymous):

Given the vector of AB as shown, draw a vector whose magnitude is half that of AB and in the opposite direction.

OpenStudy (anonymous):



OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

My answer is AB (arrow above) = -1/2CD (arrow above) but the back of my textbook says it's AB (arrow above) = -2CD (arrow above).. why and how?

OpenStudy (asnaseer):

your drawing is correct

OpenStudy (asnaseer):

the length of vector AB is twice the length of vector CD - agreed?

OpenStudy (anonymous):

agree

OpenStudy (asnaseer):

and it is in the opposite direction to CD - agreed?

OpenStudy (anonymous):

agree

OpenStudy (asnaseer):

therefore: AB = -2CD

OpenStudy (anonymous):

... but doesn't that look like CD is doubled? o.O

OpenStudy (asnaseer):

no - lets say CD had a length of 3, then AB would have a length of 6. therefore 6 = 2 * 3 therefore AB = -2 * CD

OpenStudy (anonymous):

I am so confused D:

OpenStudy (asnaseer):

the question states that CD is half the size of AB, so you could also write this as: CD = -0.5 * AB

OpenStudy (anonymous):

ohh.. but how about with the answer you mentioned earlier? How did you know?

OpenStudy (asnaseer):

how did I know what?

OpenStudy (asnaseer):

these two answers are equivalent:\[\bar{AB}=-2\bar{CD}\]\[\bar{CD}=-\frac{\bar{AB}}{2}\]

OpenStudy (anonymous):

about AB = 2CD and wait why is it -0.5*AB and not placed on CD?

OpenStudy (anonymous):

ohh okay so I get the CD = -AB/2 but how about the other answer?

OpenStudy (asnaseer):

if I tell you that I have a number whose value is 6 and ask you to find another number whose value is half my numbers value - what answer would you give?

OpenStudy (anonymous):

3

OpenStudy (asnaseer):

correct, so we can either say that:\[6 = 2\times3\]or:\[3=\frac{6}{2}\]the two statements are equivalent

OpenStudy (asnaseer):

another way to see this is as follows: you agree that: CD = -AB/2 now multiply both sides by 2: 2 * CD = -AB/2 * 2 = -AB now multiply both sides by -1: -2 * CD = AB swap sides: AB = -2 * CD

OpenStudy (anonymous):

alright so how will that connect to AB = 2*CD --> AB = -2CD, CD = AB/2... oohhh!!

OpenStudy (anonymous):

thank you! :D

OpenStudy (asnaseer):

yw :)